# MS-IMAP - A Multi-Scale Graph Embedding Approach for Interpretable Manifold Learning

### v1

@@
MS-IMAP - A Multi-Scale Graph Embedding Approach for Interpretable Manifold Learning
Shay Deutsch  [^1] Corresponding author.  [^4] Emerging Capabilities Research Group, Discover Financial Services Inc., Riverwoods, IL.
shaydeu@gmail.com
   Lionel Yelibi  [^4] Emerging Capabilities Research Group, Discover Financial Services Inc., Riverwoods, IL.
ylblio001@myuct.ac.za
   Alex Tong Lin  [^4] Emerging Capabilities Research Group, Discover Financial Services Inc., Riverwoods, IL.
atlin271@gmail.com
   Arjun Ravi Kannan  [^4] Emerging Capabilities Research Group, Discover Financial Services Inc., Riverwoods, IL.
arjun.kannan@gmail.com
1 Abstract
Deriving meaningful representations from complex, high-dimensional data in unsupervised settings is crucial across diverse machine learning applications. This paper introduces a framework for multi-scale graph network embedding based on spectral graph wavelets that employs a contrastive learning approach. A significant feature of the proposed embedding is its capacity to establish a correspondence between the embedding space and the input feature space which aids in deriving feature importance of the original features. We theoretically justify our approach and demonstrate that, in Paley-Wiener spaces on combinatorial graphs, the spectral graph wavelets operator offers greater flexibility and better control over smoothness properties compared to the Laplacian operator. We validate the effectiveness of our proposed graph embedding on a variety of public datasets through a range of downstream tasks, including clustering and unsupervised feature importance.
@@ 2 Introduction
Graph Embeddings and Manifold Learning [24, 1, 27, 28, 3] are pivotal in analyzing complex data structures prevalent in diverse machine learning applications. The emergent representations from these techniques offer insights into the underlying data structure in conjunction with downstream tasks like clustering and visualization and are particularly valuable when labels are unavailable or unreliable. However, limitations of these techniques include reliance on low frequencies of the graph Laplacian and $k$-Nearest Neighbors graph connectivity, which provide limited information, including in recent contrastive learning methods such as Uniform Manifold Approximation and Projection (UMAP). To address these limitations, this paper proposes three contributions. (1) A framework employing multi-scale graph representation using contrastive learning which enhances the expressiveness of an embedding by optimizing low and high-frequency information, capturing intricate data details. (2) A characterization of the theoretical properties of the representation power of the spectral graph wavelets (SGW) operator by considering functions sampled from the Paley-Wiener spaces [22] on combinatorial graphs and a demonstration of the flexibility and better control over smoothness properties of the SGW operator in comparison to the Laplacian operator. (3) A novel sampling technique that leverages the graph structure by sampling edges or nodes from a distribution reflecting network characteristics.
@@
A key distinction between our approach and many manifold learning methods such as UMAP is that we leverage multi-scale graph representation (using SGW) to increase the dimensionality of the data thus enhancing data representation learning. On the other hand, other manifold learning techniques predominantly employ nonlinear dimensionality reduction to reduce the dimensions of the output embedding space. We use stochastic gradient descent (SGD) to optimize the embedding space and employ a novel contrastive learning approach using a 3D-tensor to capture both low and high frequencies.
@@
Another limitation of non-linear manifold learning methods is they inherently lack direct connections to input features, thereby limiting their ability to offer interpretable results which is crucial in various applications (e.g. in finance where customer behavior needs to be summarized with a few important features to design products or strategies benefiting them and in biology where easily interpretable features may lead to life saving discoveries). Although not the main focus of this article, our proposed approach establishes a mapping between the original features and the constructed optimized embeddings and enables the measurement of the importance of the original features with respect to the embedding space (see Figure.1 for an illustration of the mapping generated by the proposed embedding). To the best of our knowledge, current manifold learning techniques produce embeddings lacking explicit interpretation. In contrast, our approach demonstrates the ability to generate interpretable embedding representations, while also being competitive with the state-of-the-art graph embeddings across applications such as finance, vision, and biology.
@@ 3 Related Work
SGW serve as an efficient tool for localization in both the spectral and vertex domains, playing a crucial role in our method (see Section 5.1 for details). SGW are constructed by applying a set of bandpass filters to the eigenvalues of the graph Laplacian, enabling multi-resolution representation of graph signals. An alternative multi-scale framework for analyzing functions on graphs is Diffusion Wavelets [23]. The Geometric Graph Scattering transform and the Graph Scattering Transform [14] utilize Diffusion Wavelets to extract graph features, which are then employed in graph classification tasks. In our work, we leverage the localization properties of SGW to develop a novel feature representation scheme, achieving superior performance in our target applications.
4 Preliminaries
@@
Consider a set of points $\mathbf{x}=\left\{\mathbf{x}_{i}\right\},\,i=1,...N,\mathbf{x}_{i}\in\mathbb{R}^{D}$, which are sampled from an unknown manifold $M\subset\mathbb{R}^{D}$. An undirected, weighted graph $G=(V,\mathbf{W})$, is constructed over $\mathbf{x}$, where $V$ corresponds to the nodes and $\mathbf{W}$ to the set of edges on the graph. The adjacency matrix $\mathbf{W}=(w_{ij})$ consists of the weights $w_{ij}$ between node $i$ and node $j$. The weights can be chosen using different techniques, such as the Gaussian kernel function or adaptive graph construction. If the weights are chosen using the Gaussian kernel function then
$(\mathbf{W})_{ij}=\left\{\begin{matrix}\exp\left(\frac{-||\mathbf{x}_{i}-\mathbf{x}_{j}||_{2}^{2}}{2\sigma_{D}^{2}}\right)&\mbox{if}\>\mathbf{x}_{j}\in\text{kNN}(\mathbf{x}_{i})\\ 0&\mbox{otherwise}\end{matrix}\right.$ (1)
where is (||||)(||\,||)( | | | | ) denotes the $L2$ distance between the points $\mathbf{x}_{i},\mathbf{x}_{j}$, kNN$(\mathbf{x}_{i})$ denotes the $k$ nearest neighbors of $\mathbf{x}_{i}$, and $2\sigma_{D}^{2}$ are parameters used to construct the graph.
An alternative is to use adaptive graph construction as proposed in [21]. The global smoothness of the graph signal function ${f}\in\mathbb{R}^{N}$ (a function over the vertices of the graph $G$) is defined using the graph Laplacian quadratic form as follows:$||\bigtriangledown{f}||^{2}=\sum_{V(i,j)}{w_{ij}(f(i)-f(j)})^{2}={f}^{T}\mathbf{L}{f}$, where $\mathbf{L}$ denotes the combinatorial graph Laplacian, defined as $\mathbf{L}=\mathbf{D}-\mathbf{W}$, with $\mathbf{D}$ the diagonal degree matrix with entries $d_{ii}=d(i)$. The degree $d(i)$ of vertex $i$ is defined as the sum of weights of edges that are connected to $i$. The normalized Laplacian is defined as ${\mathbf{L}}_{N}=\mathbf{D}^{-\frac{1}{2}}\mathbf{L}\mathbf{D}^{-\frac{1}{2}}=\mathbf{I}-\mathbf{D}^{-\frac{1}{2}}\mathbf{W}\mathbf{D}^{-\frac{1}{2}}$ and its real eigenvalues are in the interval $[0,2]$. The eigenvalues and eigenvectors of $\mathbf{L}$ are $\lambda_{1},\ldots,\lambda_{N}$ and $\mathbf{\phi}_{1},\ldots,\phi_{N}$, respectively. The graph Fourier transform (GFT) $\hat{f}$ of the graph signal $f$ is defined as the expansion of $f$ in terms of the eigenvectors $\phi$ of the graph Laplacian, so that for frequency $\lambda_{l}$ we have:
$\hat{f}(\lambda_{l})=\sum_{i}{f(i)\phi_{l}^{*}(i)}.$ (2)
We will denote the matrix representations of the eigenvectors and eigenvalues of the graph Lapacian $\mathbf{L}$ as $\Phi,\lambda$, respectively. Our approach assumes that all graph signals $f_{r}\in\mathbb{R}^{N}$ used to create the embedding correspond to the input coordinates, $\mathbf{x}=\left(f_{1},f_{2},..,f_{r},..f_{D}\right)$. The output embedding space provides a multi-scale representation approximating the intrinsic manifold coordinates.
@@ 4.1 Multi-scale representations using SGW
In the last two decades, several multi-scale representations using irregular graphs have been suggested, including SGW [15] and Diffusion Wavelets [23]. In this work we focus on the multi-scale graph transform utilizing SGW. SGW provides a natural way of trading off between spectral and spatial resolution, since the SGW coefficients are localized in both domains. These wavelets are constructed using a kernel function $g(\mathbf{L})$ operator, which acts on a function by modulating each of its Fourier modes [15], that helps in capturing a trade off between vertex (spatial) and spectral localization. Spatial localization in the graph domain is implicitly controlled by a single wavelet scale parameter defined in the spectral domain, such that the more the vertex is localized in the vertex domain, the spectral band is wider. The scale parameter enables the model to adjust the effective neighborhood sizes to the properties of the data.
To represent a signal ${f}\in\mathbb{R}^{N}$ in multiple scales $S=[s_{1},s_{2},..s_{K}]$, the SGW transform is constructed as follows: Assume that $g(\lambda)$ is a filter in the spectral domain. Let $\delta_{i}\in\mathbb{R}^{N}$ be the delta function centred at the vertex $i\in G:\delta_{i}(j)=1\,\,\mbox{if}\,\,i=j$, and $\delta_{i}(j)=0$ otherwise. Given a bandpass filter $g(\lambda)$ and a wavelet centered on node $i$ at scale $s$, the wavelets $\psi(s,i),i=1,..N$ are calculated by applying them to the delta function on a single vertex $i$, given by
$\psi(s,i)=\Phi g(s\lambda)\Phi^{T}\delta_{i}$ (3)
The value of $\psi(s,i)$ with respect to a vertex $m$ can be written as $\psi(s,i)(m)=\sum_{l=1}^{N}g(s\lambda_{l})\phi_{l}^{*}(i)\phi_{l}(m)$. Given a graph signal ${f}\in\mathbb{R}^{N}$, the SGW coefficient at node $i$ and scale $s$ can be expressed as follows:
@@
Fast computation using Chebyshev polynomials: Directly computing the SGW coefficients requires calculating the entire eigensystem of the Laplacian, which is computationally intensive - $O(N^{3})$ for $N$ points. Instead, Hammond et al. [15] suggested computing the SGW using a fast algorithm based on approximating the scaled generating kernels through low-order polynomials of $\mathbf{L}$ applied to the input data (details in Appendix).
@@ 5 Our Proposed framework: Multi-scale IMAP
In this section, we introduce Multi-scale IMAP, a framework for interpretable embedding via manifold learning utilizing a multi-scale graph representation. This approach enables us to maintain global regularity and preserve local structure without sacrificing interpretability. Our method imposes a differentiable structure on the mapping $h:\mathbf{x}\rightarrow\psi_{\mathbf{x}}$, supported on a discrete graph $G=(V,\mathbf{W})$, where $\psi_{\mathbf{x}}$ represents the encoded multi-scale graph transform of $\mathbf{x}$. Multi-scale IMAP consists of three main components:
@@
In this step, our approach constructs a multi-scale graph representation by incorporating feature signals through the SGW Transform across multiple scales and graph frequencies. We introduce two methods for encoding multi-scale transform: in the first, based on simple concatenation, all of the features transformed for all scales are represented by concatenating all the SGW coefficients for each point in a vector form. In the second based on 3D tensor optimization, we construct a 3D tensor with dimensions ${K}\times{N}\times{D}$. We leverage the encoded 3D tensor structure to align the optimized embedding with all scales simultaneously to enforce differentiable structure of the transformed features.
@@
Step 3: Network features sampling strategy.
Our approach introduces a sampling scheme that selects network features based on their estimated importance in the graph network. Employed within the optimization using SGD, this adaptive strategy dynamically prioritizes and focuses on informative parts of the graph during the embedding process. This adaptability enhances the optimization process, leading to increased efficiency in generating embeddings.
5.1 Feature Representation to encode multi-scale structure in the optimized embeddings
We propose two alternative methods to encode multi-scale representations used for subsequent optimization. Note that for both method 1 and method 2, each dimension in the embedding space is constructed using a single feature in the original feature set, which is an essential characteristic that can be leveraged for interpreting graph embeddings.
@@ 5.1 Feature Representation to encode multi-scale structure in the optimized embeddings
denoting the concatenation of the vectors corresponding to the multi-scale representation $\psi_{{f}_{i}}(s_{k},:)$ and $\psi_{{f}_{j}}(s_{k},:)$. For method (i) (also noted as method 1) where all features and all scales are concatenated together, the resulting matrix $\psi_{\mathbf{x}}$ can be represented as
where $\psi_{\mathbf{x}}\in\mathbb{R}^{KD}\times\mathbb{R}^{N}$.
(ii) 3D-Tensor Encoding - Method 2 The second encoding method concatenates the multi-scale representation for all features at a fixed scale, generating a matrix representation $\psi_{\mathbf{x}}(s_{j},:,:)\in\mathbb{R}^{D}\times\mathbb{R}^{N}$ for each scale $s_{j}$. For method 2, after concatenating all $f_{i}$ at a fixed scale $s_{j}$, we have

### v4

@@
Abstract
Deriving meaningful representations from complex, high-dimensional data in unsupervised settings is crucial across diverse machine learning applications. This paper introduces a framework for multi-scale graph network embedding based on spectral graph wavelets that employs a contrastive learning approach. We theoretically show that in Paley-Wiener spaces on combinatorial graphs, the spectral graph wavelets operator provides greater flexibility and control over smoothness compared to the Laplacian operator, motivating our approach. An additional key advantage of the proposed embedding is its ability to establish a correspondence between the embedding and input feature spaces, enabling the derivation of feature importance. We validate the effectiveness of our graph embedding framework on multiple public datasets across various downstream tasks, including clustering and unsupervised feature importance.
@@ 1 Introduction
Graph Embeddings and Manifold Learning [29, 3, 34, 35, 5] are pivotal in analyzing complex data structures prevalent in diverse machine learning applications. The emergent representations from these techniques offer insights into the underlying data structure in conjunction with downstream tasks like clustering and visualization and are particularly valuable when labels are unavailable or unreliable. However, limitations of these techniques include reliance on low frequencies of the graph Laplacian and $k$-Nearest Neighbors graph connectivity, which provide limited information, including in recent contrastive learning methods such as Uniform Manifold Approximation and Projection (UMAP). To address these limitations, this paper proposes three main contributions.
(1) A framework employing multi-scale graph representation using 3D based tensor and optimization using contrastive learning which enhances the expressiveness of an embedding by optimizing low and high-frequency information, capturing intricate data details. (2) An application of manifold interpretation to feature importance allows each dimension of our proposed embedding to be uniquely traced back and associated with the original features. (3) A characterization of the theoretical properties of the representation power of the spectral graph wavelets (SGW) operator by considering functions sampled from the Paley-Wiener spaces [27] on combinatorial graphs and a demonstration of the flexibility and better control over smoothness properties of the SGW operator in comparison to the Laplacian operator.
@@
A key distinction between our approach and many manifold learning methods such as UMAP is that we leverage multi-scale graph representation (using SGW) to increase the dimensionality of the data thus enhancing data representation learning. We use stochastic gradient descent (SGD) to optimize the embedding space and employ a novel contrastive learning approach using a 3D-tensor to capture both low and high frequencies.
@@
Generally, manifold learning methods perform nonlinear dimensionality reduction by embedding the data into lower-dimensional spaces, although some approaches may embed the data into higher dimensions [16, 31]. A common limitation of these and other graph embedding methods is their lack of direct connections to the input features, which restricts their ability to produce interpretable results. This can be crucial in fields like finance, where behavioral analysis must be summarized using key features to guide decisions, or in biology, where interpretable features are essential for understanding complex systems. Our proposed approach establishes a mapping between the original features and the constructed optimized embeddings and enables the measurement of the importance of the original features with respect to the embedding space. To the best of our knowledge, current manifold learning techniques produce embeddings lacking explicit interpretation. In contrast, our approach demonstrates the ability to generate interpretable embedding representations, while also being competitive with the state-of-the-art graph embeddings across applications such as finance, vision, and biology.
@@ 2 Related Work
SGW serve as an efficient tool for localization in both the spectral and vertex domains, playing a crucial role in our method (see Section 3.1 for details). SGW are constructed by applying a set of bandpass filters to the eigenvalues of the graph Laplacian, enabling multi-resolution representation of graph signals. Also related to our approach are graph embedding and multi-scale framework using Diffusion Wavelets [28] that embed the data in higher dimensions such as the Geometric Scattering for Graph Data Analysis [17] and the Graph Scattering Transform [16] that extract graph features, which are then employed in graph classification tasks.
2.1 Preliminaries
@@
Consider a set of points $\mathbf{x}=\left\{\mathbf{x}_{i}\right\},\,i=1,...N,\mathbf{x}_{i}\in\mathbb{R}^{D}$, which are sampled from an unknown manifold $M\subset\mathbb{R}^{D}$. An undirected, weighted graph $G=(V,\mathbf{W})$, is constructed over $\mathbf{x}$, where $V$ corresponds to the nodes and $\mathbf{W}$ to the set of edges on the graph. The adjacency matrix $\mathbf{W}=(w_{ij})$ consists of the weights $w_{ij}$ between node $i$ and node $j$. The weights can be chosen using different techniques, such as the Gaussian kernel function between the points $\mathbf{x}_{i},\mathbf{x}_{j}$ using the $k$ nearest neighbors of $\mathbf{x}_{i}$, kNN$(\mathbf{x}_{i})$ (which denotes the $k$ nearest neighbors) or adaptive graph construction as proposed in [25]. $\mathbf{L}$ denotes the combinatorial graph Laplacian, defined as $\mathbf{L}=\mathbf{D}-\mathbf{W}$, with $\mathbf{D}$ the diagonal degree matrix with entries $d_{ii}=d(i)$. The degree $d(i)$ of vertex $i$ is defined as the sum of weights of edges that are connected to $i$. The normalized Laplacian is defined as ${\mathbf{L}}_{N}=\mathbf{D}^{-\frac{1}{2}}\mathbf{L}\mathbf{D}^{-\frac{1}{2}}=\mathbf{I}-\mathbf{D}^{-\frac{1}{2}}\mathbf{W}\mathbf{D}^{-\frac{1}{2}}$ and its real eigenvalues are in the interval $[0,2]$. The eigenvalues and eigenvectors of $\mathbf{L}$ are $\lambda_{1},\ldots,\lambda_{N}$ and $\mathbf{\phi}_{1},\ldots,\phi_{N}$, respectively. In this work we use the normalized Laplacian $\mathbf{L}_{N}$, which from now on we refer to as $\mathbf{L}$ for simplicity. We will denote the matrix representations of the eigenvectors and eigenvalues of the graph Lapacian $\mathbf{L}$ as $\Phi,\lambda$, respectively. Our approach assumes that all graph signals $f_{r}\in\mathbb{R}^{N}$ used to create the embedding correspond to the input coordinates, $\mathbf{x}=\left(f_{1},f_{2},..,f_{r},..f_{D}\right)$. The output embedding space provides a multi-scale representation approximating the intrinsic manifold coordinates.
@@ 2.2 Multi-scale representations using SGW
In the last two decades, several multi-scale representations using irregular graphs have been suggested, including SGW [18] and Diffusion Wavelets [28]. In this work we focus on the multi-scale graph transform utilizing SGW. SGW provides a natural way of trading off between spectral and spatial resolution, since the SGW coefficients are localized in both domains. These wavelets are constructed using a kernel function $g(\mathbf{L})$ operator, which acts on a function by modulating each of its Fourier modes [18], that helps in capturing a trade off between vertex (spatial) and spectral localization. Spatial localization in the graph domain is implicitly controlled by a single wavelet scale parameter defined in the spectral domain, such that the more the vertex is localized in the vertex domain, the spectral band is wider. To represent a signal ${f}\in\mathbb{R}^{N}$ in multiple scales $S=[s_{1},s_{2},..s_{K}]$, the SGW transform is constructed as follows: Assume that $g(\lambda)$ is a filter in the spectral domain. Let $\delta_{i}\in\mathbb{R}^{N}$ be the delta function centred at the vertex $i\in G:\delta_{i}(j)=1\,\,\mbox{if}\,\,i=j$, and $\delta_{i}(j)=0$ otherwise. Given a bandpass filter $g(\lambda)$ and a wavelet centered on node $i$ at scale $s$, the wavelets $\psi(s,i),i=1,..N$ are calculated by applying them to the delta function on a single vertex $i$, given by $\psi(s,i)=\Phi g(s\lambda)\Phi^{T}\delta_{i}$. The value of $\psi(s,i)$ with respect to a vertex $m$ can be written as $\psi(s,i)(m)=\sum_{l=1}^{N}g(s\lambda_{l})\phi_{l}^{*}(i)\phi_{l}(m)$. Given a graph signal ${f}\in\mathbb{R}^{N}$, the SGW coefficient at node $i$ and scale $s$ can be expressed as follows:
@@
Fast computation using Chebyshev polynomials: Directly computing the SGW coefficients is computationally intensive - $O(N^{3})$ for $N$ points. Instead, Hammond et al. [18] suggested computing the SGW using a fast algorithm based on approximating the scaled generating kernels through low-order polynomials of $\mathbf{L}$ applied to the input data (details in supplementary material).
@@ 3 Our Proposed framework: Multi-Scale IMAP
In this section, we introduce Multi-scale IMAP, a framework for interpretable embedding via manifold learning utilizing a multi-scale graph representation. This approach enables us to maintain global regularity and preserve local structure without sacrificing interpretability. Our method imposes a differentiable structure on the mapping $h:\mathbf{x}\rightarrow\psi_{\mathbf{x}}$, supported on a discrete graph $G=(V,\mathbf{W})$, where $\psi_{\mathbf{x}}$ represents the encoded multi-scale graph transform of $\mathbf{x}$. Multi-scale IMAP consists of two main components:
@@
In this step, our approach constructs a multi-scale graph representation by incorporating feature signals through the SGW Transform across multiple scales and graph frequencies. We suggest a 3D tensor-based approach (detailed in Algorithm 1), where we construct a 3D tensor with dimensions ${K}\times{N}\times{D}$ to provide an effective representation for encoding multi-scale transform. We leverage the encoded 3D tensor structure to align the optimized embedding with all scales simultaneously to enforce differentiable structure of the transformed features. Note that an alternative would be to use simple concatenation, all of the features transformed for all scales are represented by concatenating all the SGW coefficients for each point in a vector form. For completeness we also describe this alternative in the Appendix. The 3D tensor-based approach (detailed in Algorithm 1) provides the most effective representation for the following reasons:
(i) Joint and simultaneous optimization: using the 3D tensor-based method enables the joint and simultaneous optimization of all scales for each sample by leveraging the SGW coefficients of each spectral band within the cross-entropy loss framework while adhering to the fixed graph structure (Equation (3.2)). This approach effectively balances coarse and fine-grain resolutions, with summation performed only after the contrastive learning optimization process is complete. By smoothing both low- and high-frequency information jointly and simultaneously with respect to the graph structure, this method reduces the reliance on threshold-based techniques for removing high-frequency spectral bands, which may be less efficient. Jointly optimizing across scales ensures consistency in the optimization process for all scales with respect to a fixed pair of nodes. Although there may be correlations between overlapping SGWs corresponding to different spectral bands, each scale is associated with a distinct bandwidth of frequencies, effectively capturing graph structures at varying levels of neighborhood information. By integrating multiple scales, SGWs capture neighborhood information continuously across varying granularities, enabling a representation that strikes a balance between local and global structural details.
(ii) Enhanced Interpretability: the proposed representation establishes a one-to-one correspondence between input features and the final embedding features, making feature interpretation more straightforward and intuitive. This is demonstrated in Section 6, where we show that the embeddings produced enable effective feature importance rankings. By ensuring this direct correspondence, our approach facilitates meaningful interpretation of the embedding space. This combination of optimization efficiency and enhanced interpretability highlights the clear advantages of the 3D tensor-based representation.
@@
3.1 Step 1: Feature Representation to encode multi-scale structure
To encode multi-scale representations used for subsequent optimization we focus on a 3D tensor based method. Note that each dimension in the embedding space is constructed using a single feature in the original feature set, which is an essential characteristic that can be leveraged for interpreting graph embeddings.
3D-Tensor Encoding: The 3D-Tensor Encoding method concatenates the multi-scale representation for all features at a fixed scale, generating a matrix representation $\psi_{\mathbf{x}}(s_{j},:,:)\in\mathbb{R}^{D}\times\mathbb{R}^{N}$ for each scale $s_{j}$. After concatenating all $f_{i}$ at a fixed scale $s_{j}$, we have
@@ 3.1 Step 1: Feature Representation to encode multi-scale structure
$\mathbf{c}(\psi_{{f}_{i}}(s_{k},:,:),\psi_{{f}_{j}}(s_{k},:,:))=\psi_{{f}_{i}}(s_{k},:,:)\left|\right|\psi_{{f}_{j}}(s_{k},:,:)$